Bob:

find the average rate of change for the function f(x) = x^2 + 3x - 5 on the interval [1,5]

Bob:

@Hero @Shadow

Bob:

is it 9

Bob:

dammit

justjm:

\[\text{Average Rate of Change} = \frac{f(b)-f(a)}{b-a}\] \[\text{Average Rate of Change}_{[1,5]} = \frac{f(5)-f(1)}{5-1}\] \[= \frac{35-(-1)}{5-1}\] \[= \frac{36}{4}\] \[=9\] You are correct if you chose 9